### v1

Abstract
Context:Eccentric exoplanets offer an opportunity to study the response of an atmosphere to changing thermal forcing and the robustness of the superrotating equatorial jet seen on tidally locked hot Jupiters. However, the atmospheric dynamics on eccentric planets strongly depend on the planetary rotation period, which is difficult to constrain observationally. The ringing phenomenon, whereby the observed emission increases and decreases after the periastron passage as the flash-heated hemisphere rotates in and out of view, can provide a tight constraint on rotation.
Aims:We examine which of the highly eccentric transiting exoplanets HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b and HD 17156 b displays sufficiently strong ringing signals for the James Webb Space Telescope (JWST) to detect.
Methods:We implemented the treatment of eccentricity and non-synchronous rotation in the non-grey climate model expeRT/MITgcm and generate synthetic light curves.
Results:We find four detectable ringing peaks on HD 80606 b and some undetectable ringing on TOI-4127 b and HD 17156 b. The lack of clouds, photo-chemistry and obliquity in our models may lead us to overestimate the amplitude of the ringing however. The strength of the ringing signal is mostly determined by the eccentricity, planetary rotation period, planet-to-star radius ratio and apparent magnitude of the system. We seek more exoplanets that could show ringing but find no candidates as promising as HD 80606 b.
Conclusions:We recommend prioritising HD 80606 b as a target for ringing with JWST. A baseline of five days after the periastron passage captures three ringing peaks, sufficient to tightly constrain the planetary rotation period. An extension to seven days adds a fourth peak, verifying the rotation period.
Examining the detectability of ringing on highly eccentric exoplanets
Key Words.:
Planets and satellites: atmospheres – Planets and satellites: individual: (HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b, HD 17156 b)1 Introduction
In the past few decades there have been great advances in the study of exoplanetary atmospheres. The observations of the recently launched James Webb Space Telescope (JWST) have contributed greatly to these efforts already (e.g. JWST Transiting Exoplanet Community Early Release Science Team et al., 2023; Feinstein et al., 2023; Tsai et al., 2023a; Dyrek et al., 2024). Because of the complicated nature of exoplanetary atmospheres, observations ought to be compared to global circulation models (GCMs) to be interpreted properly. These GCMs require well constrained orbital and planetary parameters, such as the planet’s radius, rotation period and orbital eccentricity, as input Showman et al. (2020). If the planet transits its parent star most of these quantities can be reliably measured with the notable exception of the rotation period (Sing, 2018).
Most studies so far have focussed on hot Jupiters because their size and brightness makes it relatively easy to achieve a high observational signal-to-noise (Showman et al., 2020). As a result of their close-in orbit, hot Jupiter are often on a tidally circularised orbit. Based on timescale arguments it can then safely be assumed that the planet’s rotation period is equal to the orbital period (Hut, 1981; Socrates et al., 2012; Rauscher, 2017) and its obliquity is zero (Rauscher, 2017; Rauscher et al., 2023). Consequently, the irradiation these planets receive is constant in intensity and always illuminates the same hemisphere. On the dayside hemisphere, the illumination creates a pattern of geostrophic winds that transports angular momentum towards the equator (Showman & Polvani, 2010, 2011). Consequently, a remarkable circulation pattern forms dominated by an equatorial superrotating jet, an eastwards flow of gas centred on the equator extending around the planet and typically reaching speeds of several km s. Such jets are well established both theoretically (e.g. Showman & Guillot, 2002; Showman et al., 2009; Tsai et al., 2014; Amundsen et al., 2016; Carone et al., 2020; Showman et al., 2020) and observationally (e.g. Knutson et al., 2007; Snellen et al., 2010; Zellem et al., 2014; Louden & Wheatley, 2015; May et al., 2022). Because of the major role of the Coriolis force in atmospheric dynamics, the planet’s rotation period is a decisive parameter to these circulation patterns (e.g. Showman et al., 2009; Showman & Polvani, 2010; Rauscher & Kempton, 2014; Showman et al., 2015; Baeyens et al., 2021; May et al., 2022).
In the case of a planet on an eccentric orbit, it can no longer be assumed to rotate synchronously Hut (1981). In spite of the increased difficulty, studying eccentric and non-synchronously rotating exoplanets is worthwhile as it allows one to examine the response of the atmosphere’s dynamics, chemistry and cloudiness to changing thermal forcing (Kataria et al., 2013; Lewis et al., 2017; Tsai et al., 2023b). Moreover, transit spectroscopy of cool planets on a far-out, circular orbit is rarely available as the distance restricts the viewing angles from which the planet would transit. On an eccentric planet on the other hand, a transit is more likely to take place around periastron where the atmosphere still contains traces of the cooler regime around apastron that is otherwise difficult to observe. Finally, about half of the known Jupiter-like planets are eccentric, especially those with a longer orbital period due to the distance dependence of tidal circularisation (Malmberg & Davies, 2009; Dawson & Murray-Clay, 2013; Bitsch et al., 2020).
A number of previous studies have already examined the role of eccentricity and non-synchronous rotation on the atmospheric circulation of hot and warm Jupiters. Showman et al. (2009); Rauscher & Kempton (2014); Showman et al. (2015) used 3D GCMs to study the effect of non-synchronous rotation, concluding that a shorter rotation period weakens the equatorial jet. They disagree on the effect of increasing the rotation period though, as Rauscher & Kempton (2014) claim that a westward flow develops while Showman et al. (2009, 2015) find that it leads to a significantly broader and faster eastward jets. Nonetheless, these studies all show that the rotation period is of major importance to atmospheric circulation. In a similar vein Kataria et al. (2013) examined the impact of eccentricity in a theoretical scenario, pointing out that while jets are still ubiquitous they tend to be faster the greater the eccentricity is and change quite notably in the upper atmosphere over the orbit. This is in agreement with Lewis et al. (2014) who simulated the eccentric planet HAT-P-2 b and Lewis et al. (2017); Tsai et al. (2023b) who examined the extremely eccentric HD 80606 b. The latter studies show that the jet speed on HD 80606 b decreases between periastron passages. Lewis et al. (2017) also highlight the impact of the rotation period on the circulation, composition and lofting of clouds and consequently the photometric observations of highly eccentric planets.
During the periastron passage when an eccentric planet is briefly but intensely heated, most of the flash-heating is deposited unto the side facing the star. As the planet moves further from the star and cools, that heated hemisphere rotates into and out of view of an observer on Earth. Consequently, the total brightness of the star and planet combined may display a number of local maxima, which is known as ringing. Thus, by studying the system’s light curve after the periastron passage, one can constrain the planet’s rotation period as well as how efficiently the temperature is homogenised by circulation. Using a simple thermal model without consistent dynamics or radiative transfer, Cowan & Agol (2011) predicted the presence of such ringing on the exoplanets HAT-P-2 b, HD 80606 b and HD 17156 b among others. This stands in contrast to the work of Iro & Deming (2010) who, with another simple radiative model, found little ringing in their synthetic light curves of these three planets. Kataria et al. (2013) examined the phenomenon in a theoretical case with a 3D GCM coupled to radiative transfer, highlighting the importance of the viewing geometry. Ringing has not yet been observed on any exoplanet though. de Wit et al. (2016) and Lewis et al. (2017) used the lack of ringing in Spitzer observations of HD 80606 b to constrain its rotation period and infer the likely presence of clouds, demonstrating that even the non-detection of ringing holds valuable information in some systems. As these Spitzer observations had a short baseline up to two days after the periastron passage and shorter still at the longer wavelengths more suitable to observe ringing, this does not exclude the possibility that a future JWST observation with a longer baseline may detect ringing.
No recent work employing a 3D GCM with full coupling to a non-grey radiative transfer implementation has simulated multiple known eccentric exoplanets to predict their photometric signals. Therefore, it is as of yet unclear how many and which planets are likely to display ringing. In this study, we generate synthetic light curves of five highly eccentric transiting exoplanets to examine whether JWST would be able to detect any ringing. Our ultimate goal is to inform the planning of JWST observing time by identifying viable targets and determining the baseline required for a ringing observation.
2 Eccentric Exoplanet GCM
2.1 expeRT/MITgcm
We employ the atmospheric model expeRT/MITgcm (Carone et al., 2020; Schneider et al., 2022b), featuring 3D hydrodynamics coupled to a full non-grey radiative transfer solution using the pre-calculated opacity tables. It is built upon the MITgcm dynamical core of Adcroft et al. (2004) which solves the primitive equations of hydrodynamics assuming an ideal gas. The model also assumes the planet to be a rigidly rotating sphere. We use a C32 cubed-sphere grid with 47 vertical layers, the upper 41 of which are logarithmically spaced in pressure between and 100 bar and the lower six linearly spaced in steps of 100 bar down to 700 bar. A dynamical time-step of 25 s reliably obeys the Courant-Friedrichs-Lewy stability criterion (Courant et al., 1928). To smoothen small scale noise, a fourth-order Shapiro filter with a 1 day time-scale is applied (Shapiro, 1971). At the top boundary, unphysical gravity waves are dampened by a soft sponge layer, as described by Schneider et al. (2022b). Meanwhile, we apply Rayleigh friction below 200 bar to emulate magnetic drag, which Carone et al. (2020) found improves the stability of the model. We also employ the heating term due to this parameterised drag introduced by Schneider et al. (2022a).
expeRT/MITgcm features a radiative transfer treatment including scattering using the Feautrier method with accelerated lambda iteration. Following Schneider et al. (2022b), we use a radiative time-step of 100 s, meaning the radiation field is only updated after every fourth dynamical time-step. Moreover, the radiative flux is only computed for every second column and then interpolated in between. We use six correlated-k bins (Goody et al., 1989) with sixteen Gaussian quadrature points each to discretise our 0.26 – 300 m wavelength range (Schneider et al., 2022b). We use the pre-computed opacity table of the chemical equilibrium model of Mollière et al. (2019), which minimises the Gibbs free energy on a grid of 1000 temperatures at our 47 pressures. It includes the molecules HO, CO, CH, NH, CO, HS, HCN, PH and FeH and the atoms Na and K as line opacity sources, Rayleigh scattering by H and He and collisionally induced absorption by interactions between H-H, H-He and H. TiO and VO are not added in any of these models as they often produce temperature inversions in the models (e.g. Showman et al., 2009; Lewis et al., 2014; Kataria et al., 2015; Schneider et al., 2022b) which are only observed on the hottest exoplanets (Heng & Marley, 2018; Arcangeli et al., 2018; Parmentier et al., 2018; Lothringer et al., 2018). The effects of clouds and hazes are not included in our model. Since measurements of the obliquity are available for few exoplanets and typically have a large uncertainty (Rauscher, 2017; Ohno & Zhang, 2019b; Adams et al., 2019; Bryan et al., 2020, 2021), we neglect obliquity in our model. Finally, we assume each atmosphere has a solar chemical composition, as is commonly done (e.g. Showman et al., 2009; Kataria et al., 2013; Amundsen et al., 2016; Lewis et al., 2017).
2.2 Implementation of eccentricity and non-synchronous rotation
We extended expeRT/MITgcm to allow for non-synchronous rotation and eccentric orbits as it was limited to tidally locked planets before. On tidally locked planets, the incoming irradiation is constant in intensity and always centred on the same point. On an eccentric and non-synchronously rotating planet on the other hand, the irradiation varies in two ways. Firstly, the intensity changes as with distance which depends on the true anomaly as . Herein, is the semi-major axis and the eccentricity. Secondly, the position of the substellar point in the planet’s co-rotating reference frame as with the time and the rotation period.
Since there is no analytical formula to calculate for any given , we apply an iterative Newton-Raphson method to the Kepler equation (Gregory, 2006) until the error in time is smaller than of the orbital period at each radiative time-step. This method needs no more than ten iterations, making it both precise and computationally cheap. However, the radiative transfer requires more iterations when the temperature structure changes due to the changing irradiation, which can increase the total computation time of the model by a factor of up to two.
2.3 Setup of the simulations
| m | References | ||||||||||
| (au) | (d) | (…) | (d) | () | (R) | (M) | (R) | (K) | (…) | ||
| HAT-P-2 b | 0.06740 | 5.6334729 | 0.5171 | 1.8927 | 185.2 | 0.951 | 8.74 | 1.54 | 6414 | 8.69 | 1, 2, 3, 4, 5 |
| HD 80606 b | 0.4603 | 111.436765 | 0.9318 | 1.7314 | 301.213 | 1.032 | 4.1641 | 0.98 | 5645 | 9.00 | 5, 6, 7, 8 |
| TOI-3362 b | 0.153 | 18.09547 | 0.815 | 1.2997 | 51 | 1.142 | 5.03 | 1.830 | 6532 | 10.86 | 5, 9 |
| TOI-4127 b | 0.3081 | 56.39879 | 0.7471 | 6.6118 | 129.6 | 1.096 | 2.30 | 1.293 | 6096 | 11.44 | 5, 10 |
| HD 17156 b | 0.1623 | 21.216398 | 0.6768 | 3.6775 | 121.71 | 1.095 | 3.191 | 1.508 | 6079 | 8.16 | 5, 11, 12 |
(1) Southworth (2010); (2) Pál et al. (2010); (3) Triaud et al. (2010); (4) Tsantaki et al. (2014); (5) Høg et al. (2000); (6) Pearson et al. (2022); (7) Moutou et al. (2009); (8) Naef et al. (2001); (9) Dong et al. (2021); (10) Gupta et al. (2023); (11) Winn et al. (2009); (12) Nutzman et al. (2011).
In this study, we examine five exoplanets with a high eccentricity covering a range of orbital periods, including a number of targets that have been studied before and are to be revisited by JWST. The chosen planets are HAT-P-2 b (Bakos et al., 2007), HD 80606 b (Naef et al., 2001), TOI-3362 b (Dong et al., 2021), TOI-4127 b (Gupta et al., 2023) and HD 17156 b (Fischer et al., 2007). The relevant planetary, orbital and stellar parameters are listed in Table 1. The argument of periastron defines the viewing geometry, as visualised in Fig. 1. The rotation period of the planets are computed using the commonly used pseudo-synchronous period of Hut (1981) (e.g. Kataria et al., 2013; Lewis et al., 2014, 2017; Tsai et al., 2023b) given in Eq. 1, where the rotation period is synchronised with the orbital motion around the periastron passage since that is where the tidal effects are strongest. As a result, the substellar point moves little during the periastron passage.
Upon initialisation, the wind speed is zero everywhere. The temperature is set to be horizontally homogeneous with a vertical profile comprised of three parts. The upper atmosphere at pressures smaller than 1 bar is given constant temperature based on the analytical model of Parmentier et al. (2015). That model requires an estimate of the intrinsic temperature, which is computed using the relation developed in Thorngren et al. (2019). The layers deeper than 10 bar follow an adiabatic profile with a temperature of 1400 K at 1 bar, which Sainsbury-Martinez et al. (2019) found to be appropriate for the tidally locked hot Jupiter HD 209458 b. For the exoplanets modelled in this study, which have lower mean equilibrium temperatures than HD 209458 b, this deep temperature profile is likely too hot. However, Sainsbury-Martinez et al. (2019) advise using too hot a profile as atmospheric models converge more rapidly when cooling down than when heating up. Between 1 and 10 bar, the temperature profile is linearly interpolated in pressure.
To reduce the total computation time, each planet is first treated as if it were tidally locked at a distance of the semi-major axis for 1000 d to let the temperature structure partly converge. Next, it is simulated for another 1000 d as an eccentric and non-synchronously rotating planet. On HD 80606 b, TOI-3362 b and TOI-4127 b, the temperature structure or jet speed at the apastron still changed considerably between consecutive orbits, so the run continued for another 1000 d, after which the jet speed or the upper atmosphere’s temperature did not monotonically change between orbits. However, the deep atmosphere’s temperature is not fully converged in any model. Schneider et al. (2022a) simulated the ultra hot Jupiter WASP-76 b for 86000 d and even after such a long run, the deep atmosphere’s temperature had not fully converged. As such, achieving complete convergence is unfeasible. Mayne et al. (2017); Carone et al. (2020) show that the dynamics of the deep atmosphere can considerably affect the upper atmosphere. However, Tsai et al. (2023b) indicate that additional internal heat from the overly hot deep atmosphere is unlikely to strongly affect the temperature and wind of the upper atmosphere of HD 80606 b, resulting in only a minor increase in planetary emission. Since the jet speed of the deep atmosphere has converged, we do not expect the incomplete temperature convergence of the deep atmosphere to notably affect our synthetic light curves. Therefore, our models are sufficiently converged for our purposes.
2.4 Computing synthetic light curves
Synthetic light curves are calculated during the post-processing of the GCMs using prt_phasecurve222https://prt-phasecurve.readthedocs.io/en/latest/ (Schneider et al., 2022b), based on the petitRADTRANS code (Mollière et al., 2019, 2020). We first compute a detailed radiation field throughout the atmosphere, which is then interpolated to a grid of 100 points on the side of the planet facing a hypothetical observer. We then integrate the outgoing flux over those 100 points and the chosen wavelength regime, making sure to take the viewing angle into account assuming an inclination of 90°. Finally, we divide by the stellar flux in the same wavelength range. For this work, we reduced the original forty streams used in the initial radiation field computation to six streams and the wavelength resolution from 1000 to 100. This drastically reduces the computation time of the light curve while affecting the precision of a test case based on the tidally locked exoplanet HD 209458 b by less than 1%.
We generate light curves in a wavelength range of 5 – 12 m to represent the wavelength range of the Mid-InfraRed Instrument’s (MIRI) Low Resolution Spectroscopy (LRS) mode aboard JWST, which is commonly used to study exoplanets (Rieke et al., 2015; Kendrew et al., 2015), including some planets in our sample333Approved cycle 1 program #2008 A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008. To gain some understanding of how sensitive the light curve is to the wavelength regime, we also separate the wavelength range into four, approximately representing the F560W, F770W, F1000W and F1130W imaging filters of MIRI (Rieke et al., 2015).
3 Results
3.1 Pseudo-synchronously rotating models
Each of the five eccentric exoplanets in our sample features a prominent equatorial superrotating jet throughout the entire orbit. Kataria et al. (2013); Lewis et al. (2014, 2017); Tsai et al. (2023b) found a jet on highly eccentric planets as well, demonstrating that the dynamical regime is quite robust. As in these previous works, the jet in the upper atmosphere is briefly replaced by fast day-to-night wind during the periastron passage, only to be quickly restored as the planet-star distance increases again. Lewis et al. (2017) and Tsai et al. (2023b) also note that the jet speed on HD 80606 b decreases over the orbit. Examining Fig. 2, which shows the maximal jet speed of each planet over time, we see that the jet slows down from about 1400 m s to 1000 m s on HD 80606 b, in reasonable agreement with Lewis et al. (2017); Tsai et al. (2023b). TOI-4127 b’s jet similarly slows down from 1700 m s to 1400 m s. The other three planets do not undergo such a consistent deceleration because their jets are continuously driven by the fairly strong irradiation they receive, as indicated by the equilibrium temperature shown in Fig. 3. The equilibrium temperature is the black body temperature a planet needs to radiate away the irradiation. Neglecting the albedo effect, it is
Herein, and are the effective temperature and radius of the host star and is the distance to the planet. As they are not as strongly irradiated, the thermal forcing on HD 80606 b and TOI-4127 b is insufficient to continuously feed the superrotation. Instead, their jets are only driven around the periastron passage when the irradiation is intense and the substellar point does not move as much because of the assumed pseudo-synchronous rotation period, resulting in a thermal forcing regime that strongly resembles that of a tidally locked hot Jupiter. Outside the periastron passage, the jet continues under its inertia, but is slowly decelerated by friction processes.
Figure 4 displays the contribution of the planet to the total brightness of each system in different wavelength ranges. The solid black line shows the integral of the full 5 – 12 m range. The planet’s relative contribution is greater at longer wavelengths because of its cooler temperature, which is especially notable on HD 80606 b, TOI-4127 b and HD 17156 b. Each planet’s global maximum in emission occurs shortly after the periastron passage, except for HD 80606 b’s because of its illuminated side being in the observer’s view before the periastron passage, as shown in Fig. 1. After this initial periastron peak, the brightness of HAT-P-2 b and TOI-3362 b decreases mostly monotonically until the next periastron passage. TOI-4127 b, on the other hand, shows one local maximum after the first peak while HD 80606 b has four. These ringing peaks are especially high at longer wavelengths. HD 17156 b has local maxima with small amplitudes only at longer wavelengths.
3.2 Breaking pseudo-synchronicity on HD 80606 b
To demonstrate the influence of the planet’s rotation period, we repeat the HD 80606 b model with a rotation period equal to half and double the pseudo-synchronous period. Figure 5 shows how the speed of the equatorial jet differs from the pseudo-synchronous case. We note that a longer rotation period leads to a significantly faster equatorial jet. This is explained by the substellar point moving eastwards during the periastron passage if the rotation period is longer than pseudo-synchronous. This results in more irradiation getting deposited into the already hot part of the atmosphere, thereby increasing the temperature gradient and thus the geostrophic wind speed, resulting in stronger angular momentum transport to the equator and hence a faster jet (Showman et al., 2009, 2015). Moreover, the jet is broader with a longer rotation period as explained by the equatorial Rossby deformation length , a measure of the typical jet width given by
with the Brunt-Väisälä frequency, Boltzmann’s constant, the temperature, the mean molecular weight, the atomic mass and the surface gravity (Showman & Guillot, 2002). Consequently, the re-distribution of heat becomes far more efficient with a longer rotation period. When decreasing the rotation period, the opposite occurs. The substellar point rapidly moves west at all times, even during the periastron passage, so the irradiation is spread thin and the resulting weak temperature gradient fails to pump sufficient angular momentum to the equator.
Fig. 6 demonstrates the importance of the planetary rotation period to the ringing phenomenon, as the case with a longer rotation period displays only one strong peak for two reasons. Firstly, heat is distributed more effectively by the circulation, which reduces the temperature contrast between the flash-heated and unheated hemispheres and hence the ringing amplitude. Secondly, as the rotation period is slower, it takes longer for the heated side to come into view, allowing it more time to cool and for the heat to be distributed. For the opposite reasons, the case with a short rotation period shows ten local maxima, the first of which are taller than in the pseudo-synchronously rotating case. Interestingly, the rapidly rotating case still shows small peaks over 8 d after the periastron passage, whereas the pseudo-synchronous case’s last local maximum occurs about 6 d after the periastron passage, demonstrating that the lowered efficiency of heat redistribution by circulation plays an important role.
4 Discussion
4.1 Timing of ringing peaks
Due to the strong increase in heating, the emission of the planet peaks sharply around the periastron passage on highly eccentric exoplanets. The timing of this peak relative to the periastron passage is affected by a number of parameters, including the planet’s rotation period, thereby offering a rare opportunity for an observational constraint on planetary rotation, a route taken by de Wit et al. (2016) and Lewis et al. (2017). Consider as an example the emission of HAT-P-2 b, where about half of the dayside is in the viewing disk of an observer on Earth, as visualised in Fig. 1. In that case, one might expect the dayside to come entirely into view after h. However, our synthetic light curve shown in Fig. 4 predicts that the maximum in observed emission would occur h after the periastron passage. Here we ought to include the effect of the superrotating equatorial jet which pushes the gas heated during the heating flash eastwards, hence bringing it into view earlier. We may roughly estimate the jet-enhanced timing of the periastron peak to be
where is the eastward speed at the equator due to rotation and the jet speed after periastron. For HAT-P-2 b, we estimate m s from Fig. 2 and m s, resulting in h. This matches the predicted peak timing of h given the temporal resolution of about h.
The maximum in emission is also expedited by the cooling of the gas after the periastron passage as cooling leads to a decrease in emission over time. This is especially important shortly after the periastron passage, when the planet is still hot and the emission decreases more rapidly, as demonstrated by the dependence of the radiative cooling timescale given by
with the heat capacity at constant pressure and the Stefan-Boltzmann constant (Showman & Guillot, 2002). On TOI-4127 b for instance, we estimate the maximum to occur h after the periastron passage using Eq. (4), while the synthetic light curve predicts h after the periastron passage. As such, the cooling after periastron passage and the degeneracy between rotation and the superrotating jet seriously complicate the extraction of the rotation period from the periastron peak, leading to large uncertainties in studies such as de Wit et al. (2016).
Under the right circumstances, ringing can isolate the rotation period from the timing of the ringing peaks. In the HD 80606 b simulation, the spacings between the peaks are d, d, d and d with a resolution of about d while the rotation period is d. The spacing between the periastron peak and first and second ringing peaks is shorter than the rotation period. To account for the superrotating jet, we use
Using m s and m s for HD 80606 b, we estimate d. This underestimates the modelled peak spacings in part because of the thinness of the jet due to HD 80606 b’s short rotation period and low temperature reducing the equatorial Rossby deformation radius (see Eq. (3)). Figure 7 shows the temperature evolution on HD 80606 b after periastron passage at 1 bar as seen by an observer on Earth. Initially, the flash heated gas is concentrated in one hemisphere. As the thin equatorial jet drags the hot gas at the equator eastwards, the spacing between ringing peaks is smaller than the rotation period. However, as the jet is quite thin, some flash-heated gas remains to the north-east and south-east of the substellar point at the periastron, indicated by the green dot. After an advection timescale given by
, the temperature on the equator is mostly homogenised by the jet. On HD 80606 b we estimate d, matching the fourth panel of Fig. 7. Yet the parcels of flash-heated gas outside the jet still maintain their higher temperature. As that gas continues to rotate in and out of view the last two ringing peaks follow with a spacing consistent with the rotation period. Consequently, the degeneracy between the rotation period and wind speed is broken, so the rotation period can be directly measured from the spacing between the last peaks.
4.2 Detecting ringing on these five exoplanets
Figure 4 predicts that HD 80606 b and TOI-4127 b show some ringing in the MIRI LRS wavelength regime, while HD 17156 b would have some small local maxima at long wavelengths. For a solar-like star with an apparent magnitude of 9 such as HD 80606, MIRI is expected to reach a precision of approximately 25 ppm (Greene et al., 2016; Batalha et al., 2019). As such, we might expect at the four ringing peaks on HD 80606 b to be detectable. The first local maximum on TOI-4127 b may seem detectable as well, although its second, smaller bump would not be. However, considering its -band magnitude is 11.44, the precision lies around 75 ppm instead, which the first ringing peak does not attain. Finally, the small local maximum at long wavelengths for HD 17156 b would not be detected, especially as one would have to limit oneself to the longer wavelengths of the instrument. As such, of the five exoplanets examined in this study, only HD 80606 b would exhibit any detectable ringing.
Lewis et al. (2017) demonstrated that clouds may seriously diminish ringing on HD 80606 b. Since our model does not include the effects of clouds, we likely overestimate the amplitude of the ringing peaks. Similarly, Tsai et al. (2023b) show that including photo-chemistry in the post-processing of their HD~80606~b models slightly reduces the emission magnitude after the periastron passage. However, unlike clouds this is unlikely to qualitatively alter the light curve as the composition at photospheric altitudes returns close to pre-periastron levels by the time the first ringing peak occurs. On top of the lack of clouds, our model also assumes an obliquity of 0. HD 80606 b, TOI-4127 b and HD 17156 b have an orbital period greater than the 20 d limit beyond which one cannot safely assume tides to have aligned the rotational and orbital axes (Rauscher et al., 2023). Using a simple 2D shallow-water model, Ohno & Zhang (2019a, b) showed how the obliquity and corresponding viewing geometry can strongly affect the light curve, which was confirmed by Rauscher et al. (2023). If, for instance, one of the poles receives most of the flash-heating at the periastron, that heat would not be brought in and out of view by rotation, hence no ringing. This too points towards our synthetic light curves overestimating the amplitude of the ringing. If little or no ringing is detected on HD 80606 b, this could be explained by a rotation period longer than pseudo-synchronous, clouds or obliquity. If clouds were responsible, one might expect to detect them using transit spectroscopy. If the non-detection of ringing were due to a long rotation period or an oblique orbit, the planet would prove itself useful to the research of tidal evolution and possibly to the detection of obliquity using secondary eclipse mapping (Rauscher, 2017; Rauscher et al., 2023).
In our synthetic light curve of HD 80606 b, the final ringing peak takes place 6.5 d after the periastron passage. To include the decline after that local maximum, one would need a baseline stretching until approximately 7 d after the periastron passage. As the global maximum occurs shortly before the periastron passage, the baseline should be approximately 8 d long to cover all peaks. This baseline is considerably longer than those of currently accepted proposals for JWST observations of HD 80606 b which were allotted a prime time of approximately one day444Approved cycle 1 program #2008 A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008 and Approved cycle 1 program #2488 Real Time Exoplanet Meteorology: Direct Measurement of Cloud Dynamics on the High-Eccentricity Hot Jupiter HD80606 b. https://www.stsci.edu/jwst/science-execution/program-information?id=2488. As the amplitude of the ringing in our synthetic light curves is likely overestimated, it is uncertain whether the fourth ringing peak is indeed detectable. Therefore, it may be reasonable to shorten the baseline to omit the fourth ringing peak, in which case the baseline would stretch until approximately 5 d after the periastron passage. Alternatively, a weaker constraint can be placed on the rotation period through a comparison of observations to GCMs with different rotation periods as in Fig. 6. To clearly distinguish the case from the case of HD 80606 b, a run ought to last for about 3 d after the periastron when the observed emission on the case increases while the emission continues to decrease.
4.3 Why only HD 80606 b displays ringing
In our sample of five highly eccentric exoplanets, no planet displays any detectable ringing peaks, except for HD 80606 b which has four. This remarkable difference in the level of ringing may be explained by a number of factors. Firstly, the extremely eccentric orbit of HD 80606 b results in a great difference in irradiation at the periastron compared to throughout most of the orbit, resulting in a large temperature contrast. As ringing is the consequence of a difference in emission between two hemispheres, a high eccentricity increases the amplitude of ringing peaks. Furthermore, under the assumption that the planet rotates pseudo-synchronously, a high eccentricity leads to a short rotation period, meaning little of the heat is distributed by circulation or irradiated away between rotations. Indeed, HD 80606 b has the second shortest pseudo-synchronous rotation period in our sample (see Table 1). However, Kataria et al. (2013) found an increase in the speed of the equatorial jet with eccentricity, which can raise the effectiveness of the circulation. Nevertheless, their work still suggests that a high eccentricity is beneficial for ringing, as they found no ringing on a hypothetical planet with and strong ringing on a similar case with .
Secondly, ringing in our sample is anti-correlated with the mean equilibrium temperature shown in Fig. 3. HD 80606 b has the lowest mean equilibrium temperature, while the two most irradiated cases, HAT-P-2 b and TOI-3362 b do not display any ringing. This is in part due to the long radiative cooling timescale of cool gas, which helps to maintain the temperature contrast between the two hemispheres. Furthermore, weaker irradiation is known to lead to a slower equatorial jet and hence less efficient heat redistribution (Kataria et al., 2013; Showman et al., 2015). However, Kataria et al. (2013) noted that a shorter orbital period and hence a greater equilibrium temperature leads to more and higher ringing peaks in their model with . They indicate that the strengthened ringing is aided by the shortened pseudo-synchronous rotation period. As such, the net effect of the orbital period and mean equilibrium temperature is still unclear.
Thirdly, HD 80606 b has the largest ratio of planetary radius to stellar radius of our five cases. This large ratio increases the planet’s contribution to the total brightness of the system, hence increasing the ringing signal. Moreover, a large radius also means that circulation needs to travel longer to distribute heat across the entire planet, as demonstrated by the advection timescale in Eq. 7. Finally, the fairly low apparent visual magnitude of HD 80606 b means MIRI can achieve a high precision, making ringing easier to detect.
4.4 More potential ringing targets
We conducted a search through the Encyclopaedia of Exoplanetary Systems555https://exoplanet.eu/catalog/ (Schneider et al., 2011) for more ringing candidates with a high eccentricity, short pseudo-synchronous rotation period, large planet-to-star radius fraction and low visual magnitude. Figure 8 shows the observed eccentricity and orbital period of each confirmed exoplanet in the Encyclopaedia with an entry for these two quantities. Some exoplanets are not included in the graph because their eccentricity or orbital period is not given in the catalogue, even if they are available elsewhere in the literature as is the case for TOI-4127 b for example. In order to prioritise planets with well constrained observed parameters, we limit ourselves to transiting systems.
Some planets worth considering in future theoretical or observational studies of ringing are Kepler-1656 b (Brady et al., 2018), TOI-4562 b (Heitzmann et al., 2023) and Kepler-419 b (Dawson et al., 2012), highlighted by the purple stars in Fig. 8. Unfortunately, none of these planets seem as promising a target as HD 80606 b. Kepler-1656 b has a small radius of 0.45 and long pseudo-synchronous period while Kepler-419 b and TOI-4562 b orbit quite large and hot stars. Moreover, all three systems are quite dim with visual magnitudes of 11.67 for Kepler-1656 (Høg et al., 2000), 12.09 for TOI-4562 (Zacharias et al., 2013) and 13.04 for Kepler-419 (Zacharias et al., 2013). Therefore, we recommend to prioritise searching for ringing on HD 80606 b first.
5 Conclusions
We examined which eccentric transiting exoplanets are likely to display detectable ringing, a series of local maxima in their light curves due to the side that received the heating flash at periastron rotating in and out of view. To do so, we used the 3D global circulation model with full radiative transfer coupling expeRT/MITgcm and adapted it to include non-synchronous rotation and eccentric orbits. We simulated five highly eccentric transiting exoplanets, namely HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b and HD 17156 b, with the pseudo-synchronous rotation period of Hut (1981). From these atmospheric models, we generated synthetic light curves in the wavelength range of the Low Resolution Spectroscopy mode of the Mid-InfraRed Instrument aboard JWST.
We found that all five planets in our sample feature an equatorial jet throughout their entire orbit, showing that this dynamical regime is quite robust to changing irradiation due to eccentricity. However, the jets of HD 80606 b and TOI-4127 b, the planets with the longest orbital period in our sample, are only driven during the periastron passage and continue under their inertia while slowly decelerating. The planetary rotation period plays a critical role in the dynamical regime, which is exemplified well by the lack of a consistent equatorial jet when HD 80606 b’s rotation period is decreased to half pseudo-synchronous. The rotation period is also crucial to the ringing phenomenon, as it strongly affects both the amplitude and especially the number of ringing peaks. As a result, an observation of ringing or lack thereof can provide a strong constraint on the rotation period, which is otherwise difficult to get. This would allow one to test the assumption of pseudo-synchronous rotation on highly eccentric planets.
Our synthetic light curves predict local maxima on HD 80606 b and TOI-4127 b, while the other three planets only show some small humps at in the 5 – 12 m wavelength range. TOI-4127 b’s one ringing maximum is not detectable because of its rather dim parent star. For a 9 magnitude star such as HD 80606, JWST’s detection threshold is roughly 25 ppm, which is attained by four ringing peaks on HD 80606 b, making it a promising observation target. However, our models neglect the effects of clouds and obliquity, both of which can obstruct ringing. Therefore, our synthetic light curves likely overestimate the amplitude of the ringing peaks. As a result, it is uncertain how many of the four predicted peaks on HD 80606 b are truly detectable. The timing of these peaks is affected by a complicated interplay between the planetary rotation period, equatorial jet speed, viewing geometry and cooling after the periastron passage. If the ringing lasts longer than an advection timescale however, the spacing between peaks approaches the rotation period, allowing one to extract the rotation period more reliably. An observation run with a baseline stretching seven days after the periastron passage would include all predicted peaks. A five day long observation would still cover three likely detectable peaks, enough to get the spacing between peaks after an advection timescale.
From the study of these five exoplanets, we tentatively deduce what factors facilitate observational ringing. These are a high eccentricity, a low apparent magnitude of the home star, a large planet-to-star radius ratio and a short pseudo-synchronous rotation period. In our sample of five exoplanets the ringing is more prominent on planets with a lower equilibrium temperature and long orbital period, although this goes against previous works, so the role of the orbital period is still unclear. Based on these conditions, we searched for other transiting exoplanets in the Encyclopaedia of Exoplanetary Systems that may display observable ringing, but did not find a candidate as promising as HD 80606 b. Therefore, we conclude that HD 80606 b is currently the best known transiting exoplanet candidate for ringing.
Acknowledgements.
The research presented in this paper was part of MV’s master thesis, supervised by LD and co-supervised by RB, ADS and LC. ADS, LD and LC acknowledge funding from the European Union H2020-MSCA-ITN-2019 under Grant no. 860470 (CHAMELEON). LD further acknowledges support from the KU Leuven IDN/19/028 grant ESCHER. RB acknowledges support from the ’Origins’ investment incentive. The authors would like to thank the unknown referee for their helpful comments. The data presented in this work were collected using the ‘pleiads’ computer cluster of the Institute of Astronomy, KU Leuven, maintained by Maarten Dirickx. For this work we have made use of the GCM post-processing library gcm_toolkit Schneider et al. (2022b).References
- Adams et al. (2019) Adams, A. D., Millholland, S., & Laughlin, G. P. 2019, AJ, 158, 108
- Adcroft et al. (2004) Adcroft, A., Campin, J.-M., Hill, C., & Marshall, J. 2004, Monthly Weather Review, 132, 2845
- Amundsen et al. (2016) Amundsen, D. S., Mayne, N. J., Baraffe, I., et al. 2016, A&A, 595, A36
- Arcangeli et al. (2018) Arcangeli, J., Désert, J.-M., Line, M. R., et al. 2018, ApJ, 855, L30
- Baeyens et al. (2021) Baeyens, R., Decin, L., Carone, L., et al. 2021, MNRAS, 505, 5603
- Bakos et al. (2007) Bakos, G. Á., Kovács, G., Torres, G., et al. 2007, ApJ, 670, 826
- Batalha et al. (2019) Batalha, N. E., Lewis, T., Fortney, J. J., et al. 2019, ApJ, 885, L25
- Bitsch et al. (2020) Bitsch, B., Trifonov, T., & Izidoro, A. 2020, A&A, 643, A66
- Brady et al. (2018) Brady, M. T., Petigura, E. A., Knutson, H. A., et al. 2018, AJ, 156, 147
- Bryan et al. (2020) Bryan, M. L., Chiang, E., Bowler, B. P., et al. 2020, AJ, 159, 181
- Bryan et al. (2021) Bryan, M. L., Chiang, E., Morley, C. V., Mace, G. N., & Bowler, B. P. 2021, AJ, 162, 217
- Carone et al. (2020) Carone, L., Baeyens, R., Mollière, P., et al. 2020, MNRAS, 496, 3582
- Courant et al. (1928) Courant, R., Friedrichs, K., & Lewy, H. 1928, Mathematische Annalen, 100, 32
- Cowan & Agol (2011) Cowan, N. B. & Agol, E. 2011, ApJ, 726, 82
- Dawson et al. (2012) Dawson, R. I., Johnson, J. A., Morton, T. D., et al. 2012, ApJ, 761, 163
- Dawson & Murray-Clay (2013) Dawson, R. I. & Murray-Clay, R. A. 2013, ApJ, 767, L24
- de Wit et al. (2016) de Wit, J., Lewis, N. K., Langton, J., et al. 2016, ApJ, 820, L33
- Dong et al. (2021) Dong, J., Huang, C. X., Zhou, G., et al. 2021, ApJ, 920, L16
- Dyrek et al. (2024) Dyrek, A., Min, M., Decin, L., et al. 2024, Nature, 625, 51
- Feinstein et al. (2023) Feinstein, A. D., Radica, M., Welbanks, L., et al. 2023, Nature, 614, 670
- Fischer et al. (2007) Fischer, D. A., Vogt, S. S., Marcy, G. W., et al. 2007, ApJ, 669, 1336
- Goody et al. (1989) Goody, R., West, R., Chen, L., & Crisp, D. 1989, J. Quant. Spec. Radiat. Transf., 42, 539
- Greene et al. (2016) Greene, T. P., Line, M. R., Montero, C., et al. 2016, ApJ, 817, 17
- Gregory (2006) Gregory, R. D. 2006, Classical Mechanics (Cambridge University Press)
- Gupta et al. (2023) Gupta, A. F., Jackson, J. M., Hebrard, G., et al. 2023, arXiv e-prints, arXiv:2303.14570
- Heitzmann et al. (2023) Heitzmann, A., Zhou, G., Quinn, S. N., et al. 2023, AJ, 165, 121
- Heng & Marley (2018) Heng, K. & Marley, M. S. 2018, in Handbook of Exoplanets, ed. H. J. Deeg & J. A. Belmonte (Springer), 102
- Høg et al. (2000) Høg, E., Fabricius, C., Makarov, V. V., et al. 2000, A&A, 355, L27
- Hut (1981) Hut, P. 1981, A&A, 99, 126
- Iro & Deming (2010) Iro, N. & Deming, L. D. 2010, ApJ, 712, 218
- JWST Transiting Exoplanet Community Early Release Science Team et al. (2023) JWST Transiting Exoplanet Community Early Release Science Team, Ahrer, E.-M., Alderson, L., et al. 2023, Nature, 614, 649
- Kataria et al. (2015) Kataria, T., Showman, A. P., Fortney, J. J., et al. 2015, ApJ, 801, 86
- Kataria et al. (2013) Kataria, T., Showman, A. P., Lewis, N. K., et al. 2013, The Astrophysical Journal, 767, 76
- Kendrew et al. (2015) Kendrew, S., Scheithauer, S., Bouchet, P., et al. 2015, PASP, 127, 623
- Knutson et al. (2007) Knutson, H. A., Charbonneau, D., Allen, L. E., et al. 2007, Nature, 447, 183
- Lewis et al. (2017) Lewis, N. K., Parmentier, V., Kataria, T., et al. 2017, arXiv e-prints, arXiv:1706.00466
- Lewis et al. (2014) Lewis, N. K., Showman, A. P., Fortney, J. J., Knutson, H. A., & Marley, M. S. 2014, ApJ, 795, 150
- Lothringer et al. (2018) Lothringer, J. D., Barman, T., & Koskinen, T. 2018, ApJ, 866, 27
- Louden & Wheatley (2015) Louden, T. & Wheatley, P. J. 2015, ApJ, 814, L24
- Malmberg & Davies (2009) Malmberg, D. & Davies, M. B. 2009, MNRAS, 394, L26
- May et al. (2022) May, E. M., Stevenson, K. B., Bean, J. L., et al. 2022, AJ, 163, 256
- Mayne et al. (2017) Mayne, N. J., Debras, F., Baraffe, I., et al. 2017, A&A, 604, A79
- Mollière et al. (2020) Mollière, P., Stolker, T., Lacour, S., et al. 2020, A&A, 640, A131
- Mollière et al. (2019) Mollière, P., Wardenier, J. P., van Boekel, R., et al. 2019, A&A, 627, A67
- Moutou et al. (2009) Moutou, C., Hébrard, G., Bouchy, F., et al. 2009, A&A, 498, L5
- Naef et al. (2001) Naef, D., Latham, D. W., Mayor, M., et al. 2001, A&A, 375, L27
- Nutzman et al. (2011) Nutzman, P., Gilliland, R. L., McCullough, P. R., et al. 2011, ApJ, 726, 3
- Ohno & Zhang (2019a) Ohno, K. & Zhang, X. 2019a, ApJ, 874, 1
- Ohno & Zhang (2019b) Ohno, K. & Zhang, X. 2019b, ApJ, 874, 2
- Pál et al. (2010) Pál, A., Bakos, G. Á., Torres, G., et al. 2010, MNRAS, 401, 2665
- Parmentier et al. (2015) Parmentier, V., Guillot, T., Fortney, J. J., & Marley, M. S. 2015, A&A, 574, A35
- Parmentier et al. (2018) Parmentier, V., Line, M. R., Bean, J. L., et al. 2018, A&A, 617, A110
- Pearson et al. (2022) Pearson, K. A., Beichman, C., Fulton, B. J., et al. 2022, AJ, 164, 178
- Rauscher (2017) Rauscher, E. 2017, ApJ, 846, 69
- Rauscher et al. (2023) Rauscher, E., Cowan, N. B., & Luger, R. 2023, AJ, 165, 261
- Rauscher & Kempton (2014) Rauscher, E. & Kempton, E. M. R. 2014, ApJ, 790, 79
- Rieke et al. (2015) Rieke, G. H., Wright, G. S., Böker, T., et al. 2015, PASP, 127, 584
- Sainsbury-Martinez et al. (2019) Sainsbury-Martinez, F., Wang, P., Fromang, S., et al. 2019, A&A, 632, A114
- Schneider et al. (2022a) Schneider, A. D., Carone, L., Decin, L., Jørgensen, U. G., & Helling, C. 2022a, A&A, 666, L11
- Schneider et al. (2022b) Schneider, A. D., Carone, L., Decin, L., et al. 2022b, A&A, 664, A56
- Schneider et al. (2011) Schneider, J., Dedieu, C., Le Sidaner, P., Savalle, R., & Zolotukhin, I. 2011, A&A, 532, A79
- Shapiro (1971) Shapiro, R. 1971, Journal of Atmospheric Sciences, 28, 523
- Showman et al. (2009) Showman, A. P., Fortney, J. J., Lian, Y., et al. 2009, ApJ, 699, 564
- Showman & Guillot (2002) Showman, A. P. & Guillot, T. 2002, A&A, 385, 166
- Showman et al. (2015) Showman, A. P., Lewis, N. K., & Fortney, J. J. 2015, ApJ, 801, 95
- Showman & Polvani (2010) Showman, A. P. & Polvani, L. M. 2010, Geochim. Res. Lett., 37, L18811
- Showman & Polvani (2011) Showman, A. P. & Polvani, L. M. 2011, ApJ, 738, 71
- Showman et al. (2020) Showman, A. P., Tan, X., & Parmentier, V. 2020, Space Sci. Rev., 216, 139
- Sing (2018) Sing, D. K. 2018, arXiv e-prints, arXiv:1804.07357
- Snellen et al. (2010) Snellen, I. A. G., de Kok, R. J., de Mooij, E. J. W., & Albrecht, S. 2010, Nature, 465, 1049
- Socrates et al. (2012) Socrates, A., Katz, B., & Dong, S. 2012, arXiv e-prints, arXiv:1209.5724
- Southworth (2010) Southworth, J. 2010, MNRAS, 408, 1689
- Thorngren et al. (2019) Thorngren, D., Gao, P., & Fortney, J. J. 2019, ApJ, 884, L6
- Triaud et al. (2010) Triaud, A. H. M. J., Collier Cameron, A., Queloz, D., et al. 2010, A&A, 524, A25
- Tsai et al. (2014) Tsai, S.-M., Dobbs-Dixon, I., & Gu, P.-G. 2014, ApJ, 793, 141
- Tsai et al. (2023a) Tsai, S.-M., Lee, E. K. H., Powell, D., et al. 2023a, Nature, 617, 483
- Tsai et al. (2023b) Tsai, S.-M., Steinrueck, M., Parmentier, V., Lewis, N., & Pierrehumbert, R. 2023b, MNRAS, 520, 3867
- Tsantaki et al. (2014) Tsantaki, M., Sousa, S. G., Santos, N. C., et al. 2014, A&A, 570, A80
- Winn et al. (2009) Winn, J. N., Holman, M. J., Henry, G. W., et al. 2009, ApJ, 693, 794
- Zacharias et al. (2013) Zacharias, N., Finch, C. T., Girard, T. M., et al. 2013, AJ, 145, 44
- Zellem et al. (2014) Zellem, R. T., Lewis, N. K., Knutson, H. A., et al. 2014, ApJ, 790, 53